OpenStudy (tacobell518):

MEDAL AND FAN If the midpoint between (18, y) and (20, -15) is (19, -5), find the value of y.

OpenStudy (sophiesky):

You're basically cross multiplying I believe. \[\frac{ x }{ y }\]

OpenStudy (sophiesky):

So \[\frac{ 18 }{ y } + \frac{ 20 }{ -15 } = \frac{ 19 }{ -5 }\]

OpenStudy (sophiesky):

Right?

OpenStudy (chandowd):

Well the midpoint formula is\[[(x_1 + x_2)/2 , (y_1 + y_2)/2]\] And you have both X's but only one Y. Yet you also have the final answer.

OpenStudy (tacobell518):

yeah your right

OpenStudy (chandowd):

Yes your right

OpenStudy (sophiesky):

Cool :) So now you just solve for y.

OpenStudy (chandowd):

Right so cross multiply \[18 \times - 15\] Then take that answer an divide it by 20

OpenStudy (tacobell518):

man I'm to slow

OpenStudy (tacobell518):

-270 /20 is -13.5?

OpenStudy (tacobell518):

is that the anwser

OpenStudy (chandowd):

no not quite... think of it this way. The difference between the X's is one

OpenStudy (mathmale):

You already have the x-coordinate of the midpoint. Focus on the y-coordinate. \[y _{midpoint}=\frac{ (unknown~y)+(-15) }{ 2 }=-5\]

OpenStudy (mathmale):

solve this for "unknown y."

OpenStudy (tacobell518):

I'm confused:/

OpenStudy (mathmale):

No multiplication here. And don't focus on the midpoint's x-coordinate further; it is -5. What do you understand and what do you not understand?

OpenStudy (tacobell518):

this whole finding the mid point is new to me

OpenStudy (mathmale):

Solve this:\[\frac{ y-15 }{ 2 }=-5\]

OpenStudy (tacobell518):

that's negative 3 right

OpenStudy (tacobell518):

no its not I don't know how I got that

OpenStudy (mathmale):

Please look up "midpoint." I'm sure you'll find the appropriate formulas. For the y-coordinate of the midpoint, y ou're working with \[\frac{ y _{1}+y _{2} }{ 2 }=y _{midpoint}\]

OpenStudy (mathmale):

In this case you already know that the y-coord. of the midpoint is -5. So, solve \[\frac{ y _{endpoint}-15 }{ 2 }=-5\]

OpenStudy (mathmale):

Start by mult. both sides of this equation by 2. This will eliminate the fraction.

OpenStudy (tacobell518):

ok so -30 =-10

OpenStudy (mathmale):

-30=-10 is not a true equation. Mind trying again? You have ONE unknown; it is y. y+(-15) ------- = -5 2 Multiply (-5) by 2; you get -10. Multiply y - 15 ----- by 2. What do you get? The fraction should disappear. 2